# Prompt Optimization via Retrieved Reasoning Assets and Multi-Agent Analysis

### v1

@@ 5.3.3. Summary of Human Evaluation
The human evaluation results highlight two complementary aspects of MA-SAPO. First, the multi-agent reasoning framework generates responses that are more useful, accurate, and consistent than those from a single-agent baseline, confirming the advantage of role-specialized reasoning. Second, MA-SAPO maintains strong semantic stability during prompt optimization, ensuring that improved clarity and structure do not distort the original task intent. Together, these findings demonstrate that MA-SAPO achieves both reasoning quality and semantic controllability, reinforcing the quantitative evidence presented in Section 5.3.
5.4. Cost and Latency
Table 5. Cost and Latency Comparison. MA-SAPO requires far fewer tokens and calls than other competitive multi-agent baselines, while maintaining superior optimization performance. Lower is better ($\downarrow$).
Method # Input Tokens $\downarrow$ # Output Tokens $\downarrow$ # Calls $\downarrow$
MAD 5,217 2,170 8
MARS 17,149 6,789 60
MA-SAPO (Ours) 4,968 526 2
To complement the effectiveness results, we also compare the computational cost of different multi-agent frameworks. Table 5 reports the average number of input tokens, output tokens, and calls per instance. We observe that existing multi-agent systems such as MAD and MARS incur substantial overhead. MAD requires around 5K input tokens, 2K output tokens, and 8 calls on average due to repeated multi-round debates. MARS is even heavier, consuming more than 17K input tokens, nearly 7K output tokens, and about 60 calls per instance, reflecting its iterative planner–critic–student workflow. In contrast, MA-SAPO is significantly more efficient: by constructing reasoning assets offline and adopting a single Analyzer to Refiner loop at test time, it reduces the runtime cost to roughly 5K input tokens, 0.5K output tokens, and only 2 calls per instance.
These results highlight that MA-SAPO not only improves optimization quality but also achieves a more favorable cost–latency trade-off. While MAD and MARS achieve moderate gains at the expense of heavy resource usage, MA-SAPO provides consistent improvements while keeping computational demands.
Figure 4. Case study illustrating MA-SAPO’s multi-agent prompt optimization process. Starting from an initial user prompt and evaluation scores, the framework retrieves reasoning assets, analyzes weaknesses, and iteratively refines the prompt. The regenerated response shows improved clarity, structure, and contextual depth compared to the original.
5.5. Case Study
To further illustrate the effectiveness and practical utility of MA-SAPO, we present a case study drawn from the HelpSteer2 validation dataset. [^7] Additional case studies demonstrating diverse application scenarios are available in our code repository. This example highlights MA-SAPO’s strengths in collaborative reasoning, context-aware prompt refinement, and robust adaptation across diverse input types, thereby providing concrete evidence of the framework’s ability to enhance prompt optimization in complex reasoning scenarios.
The case shown in Figure 4 demonstrates the end-to-end prompt optimization process. It begins with an initial input prompt, “What are 10 primary ways ChatGPT will change 4th–12th grade education in the United States?”, alongside the corresponding model response and evaluation scores. MA-SAPO then dynamically retrieves three semantically and thematically relevant exemplars from the training corpus (e.g., “applying ChatGPT in language learning and teaching”). These retrieved assets introduce diverse educational contexts and prompt formulations, implicitly guiding refinements such as narrowing task scope and aligning conceptual goals with actionable applications.
The Analyzer Agent identifies weaknesses in the original prompt notably, vague task specification and overly broad phrasing-both of which risk eliciting generic responses. Building on this analysis, the Refiner Agent reformulates the prompt to preserve its educational intent while incorporating a more actionable subtask, such as exploring business opportunities tied to ChatGPT-based assessment practices. The optimized prompt exhibits clearer structure, explicit segmentation, and richer contextual depth. When regenerated with the refined prompt, the model produces a response that is more coherent, pedagogically grounded, and practically insightful.
Overall, this case underscores how MA-SAPO integrates multi-agent reasoning and retrieved exemplars to achieve superior prompt optimization, outperforming conventional single-agent or static fine-tuning approaches in terms of coherence, interpretability, and alignment with user intent.
6. Conclusion
We propose MA-SAPO, a novel multi-agent framework for interpretable and controllable prompt optimization. Compared to prior methods that either perform one-shot rewrites, rely on unguided retrieval, or employ heavy multi-agent debates, MA-SAPO explicitly transforms evaluation outcomes into reusable reasoning assets and grounds refinements in evidence-backed Analyzer to Refiner interactions. Our contributions extend beyond framework design to include comprehensive experiments on two benchmarks and multiple backbones, ablation studies on retrieval depth and agent configuration, and human evaluations of reasoning quality and semantic consistency. Extensive results demonstrate that MA-SAPO consistently outperforms existing approaches, achieving higher optimization quality while remaining efficient in cost and latency.
Despite these advantages, MA-SAPO currently depends on the quality of human-annotated scores and a fixed semi-structured schema for asset construction, while also relying on a sparse BM25 retriever and a single reward model-factors that may overlook semantically relevant exemplars and introduce evaluator bias. In future work, we plan to improve robustness by integrating a dedicated Feedback Agent to validate asset quality and mitigate bias, as well as by developing hybrid dense–sparse retrieval strategies. We also aim to extend the framework to multi-turn prompts to enable more comprehensive evaluation.

### v2

@@ F.3 Summary of human evaluation
The human evaluation results highlight two complementary aspects of MA-SAPO. First, the multi-agent reasoning framework generates responses that are more useful, accurate, and consistent than those from a single-agent baseline, confirming the advantage of role-specialized reasoning. Second, MA-SAPO maintains strong semantic stability during prompt optimization, ensuring that improved clarity and structure do not distort the original task intent. Together, these findings demonstrate that MA-SAPO achieves both reasoning quality and semantic controllability, reinforcing the quantitative evidence presented in the main text.
Appendix G Implications
This work shows that prompt optimization can move beyond black-box score chasing by explicitly converting evaluation outcomes into reusable reasoning assets (explanations, diagnoses, and edit directives) that support targeted, evidence-backed edits. In practice, MA-SAPO offers a more interpretable and controllable workflow: rather than relying on trial-and-error rewriting, practitioners can trace changes to retrieved exemplars and their metric-linked rationales, which is valuable when prompt updates must be reviewed, audited, or kept consistent across versions. Moreover, MA-SAPO enables cost-aware scaling by shifting most computation offline: once the reasoning-asset corpus is built, new prompts can be optimized with lightweight retrieval plus an Analyzer-to-Refiner step, making it suitable for maintaining prompt libraries in real systems (e.g., assistants, support bots, enterprise applications) while reducing expensive multi-round agent interactions. For instance, the key contribution of MA-SAPO is a contrastive mechanism that compares, verifies, and selects among multiple LLM-generated candidates, rather than relying on advanced retrieval. As a result, even with BM25 alone, our framework remains competitive while being cheaper and faster to run. Consequently, MA-SAPO turns evaluation traces into retrievable, auditable reasoning assets, enabling repeatable prompt updates with predictable cost.
Appendix H Prompt template & design
In this section, we provide the complete prompt templates used to orchestrate the collaborative reasoning of each agent within the MA-SAPO framework. As briefly discussed in the main text, these specialized instructions are categorized into two distinct phases. The training phase prompts guide the Metric Explainer, Diagnostician, and Action Synthesizer agents in generating structured reasoning assets from evaluation scores. Conversely, the test phase prompts direct the Analyzer and Refiner agents to sequentially optimize the user’s initial prompt based on the retrieved assets. The detailed role descriptions, core tasks, and strict output constraints for each agent are presented below.
H.1 Training phase prompts
Prompt template: Metric explainer agent
Role: You are an evaluation explainer. Task: Explain WHY each of the five scores (0–4) was assigned to the model response. Provide a cohesive paragraph that bridges the gap between numerical scores and actionable improvements. Inputs: [USER_PROMPT]: {prompt_text} [MODEL_RESPONSE]: {model_response} [SCORES]: {5 metrics (helpfulness, correctness, coherence, complexity, verbosity)} Strict Rules: Output must be a single paragraph in plain text. Cover metrics in order: helpfulness $\rightarrow$ correctness $\rightarrow$ coherence $\rightarrow$ complexity $\rightarrow$ verbosity. Ground every rationale specifically in the provided response evidence. Maintain a neutral and specific tone.
Table 7: Prompt template for the metric explainer agent.
Prompt template: Diagnostician agent
Role: You are a precise evaluation doctor for LLM outputs. Task: Diagnose the primary strengths and weaknesses of the response based on metric explanations. Prescribe concrete fixes to resolve identified reasoning or alignment gaps. Inputs: [USER_PROMPT]: {prompt_text} [MODEL_RESPONSE]: {model_response} [SCORES]: {5 metrics (helpfulness, correctness, coherence, complexity, verbosity)} [REASONING_PARAGRAPH]: {reasoning_paragraph} Strict Rules: Rely strictly on the prompt, response, and provided reasoning. Address each metric once with a brief, specific rationale. Flag any mismatches between the numerical scores and the qualitative reasoning. Conclude with a crisp prescription for the required response elements.
Table 8: Prompt template for the diagnostician agent.
Prompt template: Action synthesizer agent
Role: You are a suggestion provider. Task: Synthesize the diagnosis into a single actionable prompt suggestion. Focus on enhancing the prompt’s structural and contextual depth without altering the original intent. Inputs: [USER_PROMPT]: {prompt_text} [MODEL_RESPONSE]: {model_response} [SCORES]: {5 metrics (helpfulness, correctness, coherence, complexity, verbosity)} [DIAGNOSIS]: {diagnosis_text} Strict Rules: Output ONLY the improvement suggestion (no conversational filler). Maintain the core topic and original objective of the user prompt. Focus on actionable directions (e.g., adding constraints, requesting specific formats). Balance the five metrics to guide the model toward a higher-quality response.
Table 9: Prompt template for the action synthesizer agent.
H.2 Test phase prompts
Prompt template: Analyzer agent
Role: You are the analyzer agent. Task: Conduct an independent evaluation of the prompt’s inherent qualities. Integrate insights from the three previous agents (Metric Explainer, Diagnostician, Action Synthesizer) into a unified analysis. Inputs: [USER_PROMPT]: {prompt} [RETRIEVED_CONTENT]: {retrieved_content} (Aggregated agent outputs) Strict Rules: Prioritize the Action Synthesizer’s output when synthesizing final improvement points. Eliminate redundant observations across agent outputs. Ensure all proposed improvements are concrete and actionable. Output the final analysis in a structured, cohesive paragraph.
Table 10: Prompt template for the analyzer agent.
Prompt template: Refiner agent
Role: You are the refiner agent. Task: Optimize the original prompt based on the Analyzer Agent’s feedback. Produce a final, high-performance prompt that maximizes clarity and task effectiveness. Inputs: [USER_PROMPT]: {prompt} [ANALYZER_FEEDBACK]: {analyzed_result} Strict Rules: Directly incorporate the prioritized improvements from the feedback. Preserve the original intent and task objective perfectly. Remove any unnecessary elaboration to ensure the prompt remains concise. Output ONLY the final optimized prompt.
Table 11: Prompt template for the refiner agent.
Appendix I Case study
To further illustrate the effectiveness and practical utility of MA-SAPO, we present three case studies drawn from the HelpSteer2 validation dataset. [^11] Additional case studies demonstrating diverse application scenarios are available in our code repository. Rather than merely reiterating the optimization pipeline, these examples highlight MA-SAPO’s versatile problem-solving capabilities across different types of prompt deficiencies: (1) structuring broadly scoped queries, (2) contextualizing highly open-ended requests, and (3) disambiguating subjective terminology. Across all scenarios, the core multi-agent pipeline remains consistent: specialized agents (e.g., the Analyzer and Refiner) collaboratively evaluate retrieved exemplars to dynamically synthesize a targeted optimization strategy. By examining these distinct scenarios, we provide concrete evidence of the framework’s ability to guide language models toward more coherent, interpretable, and intent-aligned responses.
I.1 Case 1: structuring broadly scoped tasks
Figure 8: Case 1. For a broad query on ChatGPT’s educational impact, retrieved exemplars signal how to frame actionable tasks. The system identifies missing operational boundaries and restructures the prompt to demand a specific format (a 10-point list with explanations). This structural constraint shifts the final response from generic statements to pedagogically grounded, practical insights.
I.2 Case 2: contextualizing open-ended requests
Figure 9: Case 2. For an under-specified request for “new novel ideas,” retrieved exemplars highlight the need for a thematic anchor. The framework overhauls the prompt by enforcing explicit domain constraints and multi-perspective exploration. This enriched context transforms the final response from boilerplate advice into deep, domain-specific conceptualization.
I.3 Case 3: disambiguating subjective terminology
Figure 10: Case 3. For a query asking for the “strongest” character, retrieved exemplars provide cues for multi-dimensional evaluation. The framework refines the prompt by decomposing the subjective term “strongest” into explicit criteria (physical, strategic, and magical abilities). These multi-faceted constraints yield a final answer that replaces a simple name-drop with a nuanced, criteria-driven assessment.